The cost, C, of a manufacturing and selling x units of a product is C = 23x+73, and the corresponding revenue, R, is= x^2-35. Find the break-even value of x?

A. 4
B. 31
C. 31 and 4
D. 27
1
Expert's answer
2012-04-17T09:56:35-0400
The profit is P = R - C = x^2 - 23x - 108 .
The break-even value is x, when P
= 0:
x^2 - 23x - 108 = 0
x = {27, -4}
The value is 27
Answer: D. 27